OpenStudy (anonymous):

Simplify the expression 4√6+√54

OpenStudy (anonymous):

@RoseDryer please help again? thank you so much.

OpenStudy (anonymous):

@xXKittyXx kausarsalley got this one.

OpenStudy (anonymous):

\[4\sqrt6 + \sqrt{54}\] \[4\sqrt6 + \sqrt{9 \times 6}\] \[4\sqrt6 + \sqrt9 \times \sqrt6\] \[4\sqrt6+3\sqrt6\] \[7\sqrt6\]

OpenStudy (anonymous):

7 √ 6 is the final answer then?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw..(: